OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 67. n + (n + 2) = 67 n − (n + 1) = 67 n + n = 67 n + (n + 1) = 67

jimthompson5910 (jim_thompson5910):

consecutive integers are whole numbers like 1, 2, 3, 4, ... basically numbers that follow one after the other

jimthompson5910 (jim_thompson5910):

so let's say 'n' is the first (smaller) integer

jimthompson5910 (jim_thompson5910):

what is the next consecutive integer

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

n is the first integer ____ is the next consecutive integer

OpenStudy (anonymous):

O

jimthompson5910 (jim_thompson5910):

still no

jimthompson5910 (jim_thompson5910):

say n = 3 the next integer would be 4

jimthompson5910 (jim_thompson5910):

4 = 3+1

jimthompson5910 (jim_thompson5910):

so replace 3 with what it is equal to

OpenStudy (anonymous):

so 4

jimthompson5910 (jim_thompson5910):

3 = ???

jimthompson5910 (jim_thompson5910):

look above to what i set 3 equal to

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

i said n = 3 above, see it

jimthompson5910 (jim_thompson5910):

so n = 3 4 is the next consecutive integer so 4 = 3+1 4 = n + 1 ... replace 3 with n (since n = 3)

jimthompson5910 (jim_thompson5910):

if n is the first, then n+1 is the next consecutive integer

OpenStudy (anonymous):

um im so confused

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

the whole thing

jimthompson5910 (jim_thompson5910):

do you have your notes with you? or your book with you?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

ok I recommend getting them out and going over them again

jimthompson5910 (jim_thompson5910):

look for the term "consecutive integer" and problems dealing with them

jimthompson5910 (jim_thompson5910):

there's probably an example showing how to solve problems like that